OpenStudy (anonymous):

If y varies inversely as x, find the constant of variation if y = 14 when x = 8. A. 112 B. 2 C. seven-fourths D. 7

OpenStudy (anonymous):

Inverse Variation is xy=k where k is the constant of variation. k=(8)(14)

OpenStudy (anonymous):

so 8 times 14 is how I get the answer?

OpenStudy (anonymous):

10. y = k / x

OpenStudy (anonymous):

And now Im confused...

OpenStudy (anonymous):

im sorry i got that from google just got done with math and dont feel like thinking lol i will try to explain though

OpenStudy (anonymous):

If y varies inversely as x, find the constant of variation if y = 14 when x = 8. The answer to this would be 112.

OpenStudy (anonymous):

I went on mathway and its apparently 7/4

OpenStudy (anonymous):

how did you write the prblem out

hero (hero):

You can manipulate xy=k to get y = k/x Both are equivalent to each other and represent inverse variation.

OpenStudy (anonymous):

I wrote it out the same way everyone has been telling me to.

hero (hero):

@Partycool, I assumed you would have these figured out by now.

OpenStudy (anonymous):

what exactly did u put into math way